OpenStudy (anonymous):

WILL GIVE MEDALS AND FAN! If y (Subtract) 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? A. {(0, 3), (1, 6), (2, 9)} B. {(3, 0), (6, 1), (9, 2)} C. {(1, 3), (2, 6), (3, 9)} D. {(3, 1), (6, 2), (9, 3)}

OpenStudy (anonymous):

@mathstudent55 @iambatman

OpenStudy (anonymous):

(x,y) are your points, plug in x get output y

OpenStudy (anonymous):

3=3x Lets see (0,3) 3=3(0) 3 = 0 ? does that make sense?

OpenStudy (anonymous):

Yes, it does

OpenStudy (anonymous):

So, the answer id B?

OpenStudy (anonymous):

Well think about it, look at the first option they give, (3,0) 3=3(3) => 3 = 9 that's not right. You want the output to be equal on both sides, but if you had (1,3) 3=3(1) 3=3 that's right!

OpenStudy (anonymous):

So it's D?

OpenStudy (anonymous):

Try it yourself...

OpenStudy (anonymous):

Plug in the values, see if it makes sense to you.

OpenStudy (anonymous):

No it seems like the answer is C... I think. I'm almost positive.

OpenStudy (anonymous):

Can you prove it?